OpenStudy (anonymous):

Solve the following equation: y=6x + 1 2x - 9y =30

OpenStudy (anonymous):

OK now I see, it was just my browser...

OpenStudy (anonymous):

There are 2 methods to solve this. Substitution and elimination. I prefer elimination but we can do the other if you wish. Choose! :-)

OpenStudy (anonymous):

actually you can solve it by graphing too, but I wont do that now

OpenStudy (anonymous):

OK I do the elimination there it is important to make the same amount of x-s or y-s. here we have y=6x+1 if we multiply it by 9 than we will get 9y=54x+9 the second equation is 2x - 9y =30 so we have 9y in the 1st and -9y in the 2nd so if we add them together they will drop out. So we have to add the 2 equations to get. 9y+2x-9y=54x+9+30 2x=54x+39 52x=-39 x=-39/52

OpenStudy (anonymous):

Thanks. I think i get it now. i "THINK"

OpenStudy (anonymous):

I know you!